# Intermediate mass-ratio inspirals in a dense dark-matter environment: Effects of the initial dark-matter distribution

### v1

@@ V.1 Dephasing for second-generation inspirals
Table 2: Gravitational-wave dephasing for a second-generation inspiral relative to vacuum. We denote $N_{\mathrm{cycles}}^{\mathrm{2_{II}}}$ as the number of GW cycles between a given starting radius and ISCO for a simulation with $j_{\mathrm{min}}=\sqrt{8}$ for a second-generation inspiral. The GW dephasing between a vacuum inspiral and one with an angular-momentum cutoff ($j_{\mathrm{min}}=\sqrt{8}$) for a second-generation merger is given by $\Delta N_{\mathrm{cycles}}^{\mathrm{(0-2_{II})}}$. As discussed further in Sec. IV.2, we use $m_{2}=10\,\mathrm{M}_{\odot}+\delta m_{\mathrm{2,4y}}^{\mathrm{2_{II}}}$ for the vacuum binary when we compute the dephasing. The dephasing is computed for the last four years of inspiral before the secondary reaches the ISCO for the same initial radius in vacuum. We choose $\gamma_{\mathrm{sp}}=7/3$ and $\rho_{\mathrm{sp}}=200\,\mathrm{M}_{\odot}\,\mathrm{pc}^{-3}$ for the dark-matter parameters.
@@
We label the simulations of second-generation inspirals with a cutoff by $\mathrm{2_{II}}$, so as to distinguish them from the first-generation inspirals (labeled by $\mathrm{2_{I}}$). The GW dephasing of the second-generation inspirals from corresponding vacuum systems are given in Table 2. Similar to Table 1, there is a peak in the dephasing values against vacuum at a primary mass of $3\times 10^{4}\,\mathrm{M}_{\odot}$, but the fractional dephasing is largest for $m_{1}=10^{5}\,\mathrm{M}_{\odot}$. As in Sec. IV.2, we caution that when comparing the numbers between the third column from Table 2 and the sixth column from Table 1 that the systems have different masses $m^{\mathrm{A}}_{\mathrm{2,4y}}$ and inspiral from different radii $r^{\mathrm{A}}_{\mathrm{2,4y}}$. Nevertheless, the second-generation mergers have a smaller dephasing from vacuum than the first-generation mergers do, and the dephasing is smaller in second-generation binaries with less extreme mass ratios. This trend in the dephasing is related to the fact that there is more feedback for smaller primary masses. The increased feedback leads to a greater difference in the initial data between the first- and second-generation mergers as the mass ratio becomes less extreme. However, the feedback from dynamical friction and accretion can be so large during the inspiral that the density is significantly depleted and the dark-matter effects on the inspiral are reduced (which also produces the peak in the dephasing numbers).
@@
Part of the GW dephasing for the also comes from differences in the respective masses of the secondary (which is shown in the fourth column of Table 2). The behavior of $\delta m_{\mathrm{2,4y}}^{\mathrm{2_{II}}}$ is similar to the two cases of $\delta m_{\mathrm{2,4y}}^{\mathrm{A}}$ given in Table 1, but the largest value for $\delta m_{\mathrm{2,4y}}^{\mathrm{2_{II}}}$ occurs at a mass of $m_{1}=10^{4}\,\mathrm{M}_{\odot}$. The shift in the peak value and the smaller values of $\delta m_{\mathrm{2,4y}}^{\mathrm{2_{II}}}$ at less extreme mass ratios arises because of the larger changes to the densities that occur during the first inspiral. This leaves less dark matter that can be accreted onto the secondary during the second-generation inspiral. In the cases with the more-extreme mass ratios, such as when $m_{1}=10^{5}\,\mathrm{M}_{\odot}$, the accreted mass $\delta m_{\mathrm{2,4y}}^{\mathrm{2_{II}}}$ is more similar to $\delta m_{\mathrm{2,4y}}^{\mathrm{2_{I}}}$, because of the smaller changes to the density during the first-generation inspiral. In addition, because $m_{2}$ is everywhere smaller during the second inspiral (given that we always choose $m_{2,0}=10\,\mathrm{M}_{\odot})$, radiation reaction is slightly smaller, which contributes an additional source of dephasing between the first- and second-generation inspirals. These three effects together combine to create the trends shown in Table 2. In addition, a comparison of the third column of Table 2 and the fourth of Table 1 shows that for smaller primary masses, the changes in the initial densities have a larger effect on the GW dephasing than including a nonzero angular-momentum cutoff does.
@@
Figure 4: Cycles of dephasing versus gravitational-wave frequency for different initial dark-matter distributions and angular-momentum cutoffs. The dark-matter parameters ($\rho_{\mathrm{sp}}=200\,\mathrm{M}_{\odot}/\mathrm{pc}^{3}$ and $\gamma_{\mathrm{sp}}=7/3$) are the same as in Fig. 2 for both panels and the primary mass is chose to be $m_{1}=10^{3}\,\mathrm{M}_{\odot}$ (left) and $m_{1}=10^{5}\,\mathrm{M}_{\odot}$ (right). The dotted, black lines are the GW dephasing between vacuum binaries and those in a density profile without enforcing a cut to angular momentum ($j_{\mathrm{min}}=0$), while the solid blue and dash-dotted orange correspond to the GW dephasing from vacuum for systems with a nonzero angular-momentum cutoff ($j_{\mathrm{min}}=\sqrt{8}$) for first- and second-generation inspirals, respectively. The vertical, black, dotted lines are plotted at the frequency for which the IMRI will merger in four years. More detailed discussion of the results is given in the text of Sec. V.1.
@@
We now turn to the right panel of Fig. 4 (the results for a primary mass of $m_{1}=10^{5}\,\mathrm{M}_{\odot}$). For this more massive primary, the secondary spends a much larger portion of the last four years of inspiral at radii closer to the ISCO, where the effects of the nonzero angular-momentum cutoff on the density are more significant (see Fig. 1, and note that $r_{\mathrm{2,4y}}^{\mathrm{A}}\approx 17R_{\mathrm{s}}$). At all frequencies shown, the dephasing of the second-generation inspiral from vacuum is a larger fraction of that of the first-generation inspiral for the $m_{1}=10^{5}\,\mathrm{M}_{\odot}$ than it is for the $m_{1}=10^{3}\,\mathrm{M}_{\odot}$ case. This is consistent with the fact that the DM distribution following the first-generation inspiral is much more similar to the initial distribution for this merger. The dephasing curves for the first- and second-generation inspirals converge towards each other at larger frequencies, which is indicative of a smaller difference in the densities at smaller radii.
We conclude this part, as we did in Sec. IV.2 with a comment about the choice of the initial binary separation and its effects on the dephasing during the last four years of inspiral. Given that there are some artifacts in the density at radii greater than $r_{\mathrm{2,4y}}^{\mathrm{A}}$ (as noted in Sec. IV.3) it is worth revisiting whether they might affect the results for the dephasing. We perform a similar analysis to that in Sec. IV.2, where we performed simulations starting at $r_{2}=2r_{\mathrm{2,4y}}$ and $r_{2}=4r_{\mathrm{2,4y}}$, though now for second-generation inspirals. We use the densities from first-generation simulations starting at that same radius, and we compare the dephasing to that listed in Table 2. We observed the same behavior of dephasing with the initial separation as described Sec. IV.2 (that is, a difference in dephasing of less than our numerical uncertainty when starting at $r_{2}=4r_{\mathrm{2,4y}}$, but a significant difference when starting at $r_{2}=2r_{\mathrm{2,4y}}$). We conclude that the simulations are relatively robust to a somewhat different initial separation, including second-generation inspirals.
@@ V.2 Dark-matter density following the second inspiral
Figure 5: Dark-matter density at three stages of the inspiral for a second-generation inspiral. The dark matter and binary parameters are the same as in the Fig. 3. We show the density of dark matter versus radius for several orbital separations during the inspiral of a second-generation merger. The dotted black line shows the initial density profile when $r_{2}=3r_{\mathrm{2,4y}}$, the dash-dotted orange the density for $r_{2}=r_{\mathrm{2,4y}}$, and the solid blue is the density for $r_{2}=r_{\mathrm{ISCO}}$. The light-gray dashed line is the initial density from the first-generation inspiral, which is given for comparison. More detailed discussion about the figure appears in the text of Sec. V.
@@
In Fig. 5, we plot the density at three stages of the inspiral for a second-generation inspiral. The initial parameters for the binary are the same as those used in Fig. 3, which has $m_{1}=$ $10^{3}\,\mathrm{M}_{\odot}$, $m_{2,0}=$ $10\,\mathrm{M}_{\odot}$, $\gamma_{\mathrm{sp}}=7/3$, and $\rho_{\mathrm{sp}}$ = $200\,\mathrm{M}_{\odot}/\mathrm{pc}^{3}$. The density of the first-generation inspiral when the secondary reaches the ISCO (the solid blue curve in Fig. 3) serves as the initial density for the second inspiral (it is depicted as the dotted, black curve in Fig. 5). The initial density from Fig. 3 is reproduced as the dashed gray line in Fig. 5 for comparison. The curves not contained in Fig. 3 are the dashed-dotted orange curve, which represents the density during the second-generation inspiral when $r_{2}=r_{\mathrm{2,4y}}$, and the solid blue curve, which is the corresponding density when $r_{2}=r_{\mathrm{ISCO}}$. Comparing the blue and black curves shows that the density decreases significantly with each subsequent inspiral. In particular, the peak of the density near $r=4\times 10^{-9}\,\mathrm{pc}$ is decreased by slightly less than an order of magnitude after each inspiral. This leads to an overall flatter radial density profile at small radii. A stronger flattening is predicted to occur in other scenarios as well, such as in the “annihilation plateau” for self-annihilating dark-matter models [3] or some dark-matter profiles formed from some collapse scenarios (e.g., [52]).
@@
As described in Sec. IV, it is important to start the secondary at $r_{2}=3r_{\mathrm{2,4y}}$, so that the density when $r_{2}=r_{\mathrm{2,4y}}$ is more consistent with an adiabatic quasicircular inspiral from much larger radii. The local density when $r_{2}=r_{\mathrm{2,4y}}$ (the orange star) is lower than it would have been had we the inspiral started with $r_{2}=r_{\mathrm{2,4y}}$ with the density from the past inspiral (the black, dotted curve). As we did in Sec. IV.3, we again comment that the density in Fig. 5 at radii less than $r_{\mathrm{2,4y}}$ the density is reasonably robust to the choice of the initial separation, but further out it is not. Therefore, the detailed profile at radii larger than $r_{\mathrm{2,4y}}$ is not necessarily consistent with a quasicircular inspiral beginning from a much larger initial separation.
@@ VI Conclusions
There are several directions in which this work could be extended or generalized. First, we worked to leading order in mass ratio and in post-Newtonian order for all aspects of the evolution equations for the inspiral, but higher-order calculations would be necessary for modeling these systems at the accuracy required for LISA data analysis. Second, we assumed that the secondary always follows a quasicircular orbit about the primary; depending on its formation scenario, an IMRI could have nonzero eccentricity. The effects of eccentricity have been considered in [38], though not for the initial densities considered in this work. Third, we assumed that the primary remained in the center of the dark-matter spike at all times. At leading order in the mass ratio, the center of mass of the system coincides with the primary’s position, but at subleading order, it does not. How the primary moves with respect to the center of the dark-matter distribution would also be interesting to incorporate. Following a first-generation merger, the resulting black hole (which would serve as the primary for the second-generation merger) would experience a gravitational-wave recoil (a “kick”) with a speed that is proportional to the mass ratio of the system. This could also displace the primary from the center of the dark-matter density, and depending on the relaxation time, the second-generation merger could be affected by this. Any such effects would be largest for the least-extreme mass-ratio IMRIs.
@@ Appendix A Computing the dynamical-friction feedback scattering rate
When $m>1$, it is useful to perform a reciprocal modulus transformation (see [60, §19.7]), as was done in [40]. As we discuss in Appendix B, in [40], given the definition of the analogues of $\alpha_{1}$ and $\alpha_{2}$ used when $\mathcal{J}_{\mathrm{min}}=0$, it can be shown that the conditions required of the reciprocal modulus transformation (discussed below) always hold. This transformation, when written in terms of the modulus $m$ (not $k$, as in [60, §19.7]), allows an elliptic integral of the second kind with $m>1$ to be recast in terms of elliptic integrals with $1/m<1$ of the first and second kinds as follows:

### v2

@@ V.1 Dephasing for second-generation inspirals
Table 2: Gravitational-wave dephasing for a second-generation inspiral relative to vacuum. We denote $N_{\mathrm{cycles}}^{\mathrm{2_{II}}}$ [see Eq. (51)] as the number of GW cycles between a given starting radius and ISCO for a simulation with $j_{\mathrm{min}}=\sqrt{8}$ for a second-generation inspiral. The GW dephasing between a vacuum inspiral and one with an angular-momentum cutoff ($j_{\mathrm{min}}=\sqrt{8}$) for a second-generation merger is given by $\Delta N_{\mathrm{cycles}}^{\mathrm{(0-2_{II})}}$ [see Eq. (52)]. As discussed further in Sec. IV.2, we use $m_{2}=10\,\mathrm{M}_{\odot}+\delta m_{\mathrm{2,4y}}^{\mathrm{2_{II}}}$ [see Eq. (53)] for the vacuum binary when we compute the dephasing. The dephasing is computed for the last four years of inspiral before the secondary reaches the ISCO for the same initial radius in vacuum. We choose $\gamma_{\mathrm{sp}}=7/3$ and $\rho_{\mathrm{sp}}=200\,\mathrm{M}_{\odot}\,\mathrm{pc}^{-3}$ for the dark-matter parameters.
@@
We label the simulations of second-generation inspirals with a cutoff by $\mathrm{2_{II}}$, so as to distinguish them from the first-generation inspirals (labeled by $\mathrm{2_{I}}$). The GW dephasing of the second-generation inspirals from corresponding vacuum systems are given in Table 2. Similar to Table 1, there is a peak in the dephasing values against vacuum at a primary mass of $3\times 10^{4}\,\mathrm{M}_{\odot}$, but the fractional dephasing is largest for $m_{1}=10^{5}\,\mathrm{M}_{\odot}$. As in Sec. IV.2, we caution that when comparing the numbers between the third column from Table 2 and the sixth column from Table 1 that the systems have different masses $m^{\mathrm{(A)}}_{\mathrm{2,4y}}$ and inspiral from different radii $r^{\mathrm{(A)}}_{\mathrm{2,4y}}$. Nevertheless, the second-generation mergers have a smaller dephasing from vacuum than the first-generation mergers do, and the dephasing is smaller in second-generation binaries with less extreme mass ratios. This trend in the dephasing is related to the fact that there is more feedback for smaller primary masses. The increased feedback leads to a greater difference in the initial data between the first- and second-generation mergers as the mass ratio becomes less extreme. However, the feedback from dynamical friction and accretion can be so large during the inspiral that the density is significantly depleted and the dark-matter effects on the inspiral are reduced (which also produces the peak in the dephasing numbers).
@@
Part of the GW dephasing for the also comes from differences in the respective masses of the secondary (which is shown in the fourth column of Table 2). The behavior of $\delta m_{\mathrm{2,4y}}^{\mathrm{2_{II}}}$ is similar to the two cases of $\delta m_{\mathrm{2,4y}}^{\mathrm{(A)}}$ given in Table 1, but the largest value for $\delta m_{\mathrm{2,4y}}^{\mathrm{2_{II}}}$ occurs at a mass of $m_{1}=10^{4}\,\mathrm{M}_{\odot}$. The shift in the peak value and the smaller values of $\delta m_{\mathrm{2,4y}}^{\mathrm{2_{II}}}$ at less extreme mass ratios arises because of the larger changes to the densities that occur during the first inspiral. This leaves less dark matter that can be accreted onto the secondary during the second-generation inspiral. In the cases with the more-extreme mass ratios, such as when $m_{1}=10^{5}\,\mathrm{M}_{\odot}$, the accreted mass $\delta m_{\mathrm{2,4y}}^{\mathrm{2_{II}}}$ is more similar to $\delta m_{\mathrm{2,4y}}^{\mathrm{2_{I}}}$, because of the smaller changes to the density during the first-generation inspiral. In addition, because $m_{2}$ is everywhere smaller during the second inspiral (given that we always choose $m_{2,0}=10\,\mathrm{M}_{\odot})$, radiation reaction is slightly smaller, which contributes an additional source of dephasing between the first- and second-generation inspirals. These three effects together combine to create the trends shown in Table 2. In addition, a comparison of the third column of Table 2 and the fourth of Table 1 shows that for smaller primary masses, the changes in the initial densities have a larger effect on the GW dephasing than including a nonzero angular-momentum cutoff does.
@@
Figure 4: Cycles of dephasing versus gravitational-wave frequency for different initial dark-matter distributions and angular-momentum cutoffs. The dark-matter parameters ($\rho_{\mathrm{sp}}=200\,\mathrm{M}_{\odot}/\mathrm{pc}^{3}$ and $\gamma_{\mathrm{sp}}=7/3$) are the same as in Fig. 2 for both panels and the primary mass is chosen to be $m_{1}=10^{3}\,\mathrm{M}_{\odot}$ (left) and $m_{1}=10^{5}\,\mathrm{M}_{\odot}$ (right). The dotted, black lines are the GW dephasing between vacuum binaries and those in a density profile without enforcing a cut to angular momentum ($j_{\mathrm{min}}=0$), while the solid blue and dash-dotted orange correspond to the GW dephasing from vacuum for systems with a nonzero angular-momentum cutoff ($j_{\mathrm{min}}=\sqrt{8}$) for first- and second-generation inspirals, respectively. The vertical, black, dotted lines are plotted at the frequency for which the IMRI will merger in four years. More detailed discussion of the results is given in the text of Sec. V.1.
@@
We now turn to the right panel of Fig. 4 (the results for a primary mass of $m_{1}=10^{5}\,\mathrm{M}_{\odot}$). For this more massive primary, the secondary spends a much larger portion of the last four years of inspiral at radii closer to the ISCO, where the effects of the nonzero angular-momentum cutoff on the density are more significant (see Fig. 1, and note that $r_{\mathrm{2,4y}}^{\mathrm{(A)}}\approx 17R_{\mathrm{s}}$). At all frequencies shown, the dephasing of the second-generation inspiral from vacuum is a larger fraction of that of the first-generation inspiral for the $m_{1}=10^{5}\,\mathrm{M}_{\odot}$ than it is for the $m_{1}=10^{3}\,\mathrm{M}_{\odot}$ case. This is consistent with the fact that the DM distribution following the first-generation inspiral is much more similar to the initial distribution for this merger. The dephasing curves for the first- and second-generation inspirals converge towards each other at larger frequencies, which is indicative of a smaller difference in the densities at smaller radii.
As in Sec. IV.2, we comment about the choice of the initial binary separation and its effects on the dephasing during the last four years of inspiral. Given that there are some artifacts in the density at radii greater than $r_{\mathrm{2,4y}}^{\mathrm{(A)}}$ (as noted in Sec. IV.3) it is worth revisiting whether they might affect the results for the dephasing. We perform a similar analysis to that in Sec. IV.2, where we performed simulations starting at $r_{2}=2r_{\mathrm{2,4y}}$ and $r_{2}=4r_{\mathrm{2,4y}}$, though now for second-generation inspirals. We use the densities from first-generation simulations starting at that same radius, and we compare the dephasing to that listed in Table 2. We observed the same behavior of dephasing with the initial separation as described Sec. IV.2 (that is, a difference in dephasing of less than our numerical uncertainty when starting at $r_{2}=4r_{\mathrm{2,4y}}$, but a significant difference when starting at $r_{2}=2r_{\mathrm{2,4y}}$). We conclude that the GW dephasing, including that for second-generation inspirals, is robust to the choice of the initial separation, when that separation is sufficiently large.
Note that while the dephasing curves for first- and second-generation mergers shown in Fig. 4 are distinguishable from each other, we have not attempted to determine if there exists a first-generation merger with a different initial density that could mimic the dephasing curve of the second-generation merger. It would be interesting to explore this possibility in future work.
@@ V.2 Dark-matter density following the second inspiral
Figure 5: Dark-matter density at three stages of the inspiral for a second-generation inspiral. The dark matter and binary parameters are the same as in the Fig. 3. We show the density of dark matter versus radius for several orbital separations during the inspiral of a second-generation merger. The dotted black line shows the initial density profile when $r_{2}=3r_{\mathrm{2,4y}}$, the dash-dotted orange the density for $r_{2}=r_{\mathrm{2,4y}}$, and the solid blue is the density for $r_{2}=r_{\mathrm{ISCO}}$. The light-gray dashed line is the initial density from the first-generation inspiral, which is given for comparison. More detailed discussion about the figure appears in the text of Sec. V.2.
In Fig. 5, we plot the density at three stages of the inspiral for a second-generation inspiral. The initial parameters for the binaries are the same as those used in Fig. 3, which have $m_{2,0}=$ $10\,\mathrm{M}_{\odot}$, $\gamma_{\mathrm{sp}}=7/3$, and $\rho_{\mathrm{sp}}$ = $200\,\mathrm{M}_{\odot}/\mathrm{pc}^{3}$. The panel on the left shows the $m_{1}=$ $10^{3}\,\mathrm{M}_{\odot}$ case, while the panel on the right is for $m_{1}=$ $10^{5}\,\mathrm{M}_{\odot}$. The densities of the first-generation inspirals when the secondary reaches the ISCO (the solid blue curves in Fig. 3) serve as the initial densities for the second inspirals (they are depicted as the dotted, black curves in the two panels of Fig. 5). The initial densities from Fig. 3 are reproduced as the dashed gray curves in Fig. 5 for comparison. The curves not contained in Fig. 3 are the dashed-dotted orange curves, which represent the densities during the second-generation inspiral when $r_{2}=r_{\mathrm{2,4y}}$, and the solid blue curves, which are the corresponding densities when $r_{2}=r_{\mathrm{ISCO}}$.
@@
Comparing the blue and black curves shows that the densities decrease with each subsequent inspiral, but the amount of decrease depends strongly on the primary mass. In particular, for the $m_{1}=10^{3}\,\mathrm{M}_{\odot}$ case, the peak of the density near $r\approx 4\times 10^{-9}\,\mathrm{pc}$ is decreased by slightly less than an order of magnitude after each inspiral. This leads to an overall flatter radial density profile at small radii. A stronger flattening is predicted to occur in other scenarios as well, such as in the “annihilation plateau” for self-annihilating dark-matter models [16], the shallower power law in self-interacting DM models (see, e.g., [10]), or some DM profiles formed from collapse scenarios (e.g., [20]). The change in density for the $m_{1}=$ $10^{5}\,\mathrm{M}_{\odot}$ case is less significant, but there is still a noticeable redistribution of DM particles. For example, the density decreases for $r\lesssim 5\times 10^{-7}\,\mathrm{pc}$ and increasing it at radii larger than $\sim 5\times 10^{-7}$ pc.
@@
As described in Sec. IV, it is important to start the secondary at $r_{2}=3r_{\mathrm{2,4y}}$, so that the density when $r_{2}=r_{\mathrm{2,4y}}$ is more consistent with an adiabatic quasicircular inspiral from much larger radii. The local density when $r_{2}=r_{\mathrm{2,4y}}$ (the orange star) is lower than it would have been had we the inspiral started with $r_{2}=r_{\mathrm{2,4y}}$ with the density from the past inspiral (the black, dotted curve). As we did in Sec. IV.3, we again comment that the density in Fig. 5 at radii less than $r_{\mathrm{2,4y}}$ is reasonably robust to the choice of the initial separation, but further out it is not. Therefore, the detailed profile at radii larger than $r_{\mathrm{2,4y}}$ is not necessarily consistent with a quasicircular inspiral beginning from a much larger initial separation.
@@ VI Conclusions
There are several directions in which this work could be extended or generalized. First, we worked to leading order in mass ratio and in post-Newtonian order for all aspects of the evolution equations for the inspiral, but higher-order calculations would be necessary for modeling these systems at the accuracy required for LISA data analysis. Second, we assumed that the secondary always follows a quasicircular orbit about the primary; depending on its formation scenario, an IMRI could have nonzero eccentricity. The effects of eccentricity have been considered in [43, 45], though not for the initial densities considered in this work. While Fig. 14 of [43] showed that the dephasing for eccentricities below roughly 0.7 were qualitatively similar to the circular case, it would be useful to verify that this continues to hold for initial data with a cutoff in angular momentum. Third, we assumed that the primary remained in the center of the dark-matter spike at all times. At leading order in the mass ratio, the center of mass of the system coincides with the primary’s position, but at subleading order, it does not. How the primary moves with respect to the center of the dark-matter distribution would also be interesting to incorporate. Following a first-generation merger, the resulting black hole (which would serve as the primary for the second-generation merger) would experience a gravitational-wave recoil (a “kick”) with a speed that is proportional to the mass ratio of the system. This could also displace the primary from the center of the dark-matter density, and depending on the relaxation time, the second-generation merger could be affected by this. Any such effects would be largest for the least-extreme mass-ratio IMRIs.
@@ Appendix A Computing the dynamical-friction feedback scattering rate
When $m>1$, it is useful to perform a reciprocal modulus transformation (see [29, §19.7]), as was done in [46]. As we discuss in Appendix B, in [46], given the definition of the analogues of $\alpha_{1}$ and $\alpha_{2}$ used when $\mathcal{J}_{\mathrm{min}}=0$, it can be shown that the conditions required of the reciprocal modulus transformation (discussed below) always hold. This transformation, when written in terms of the modulus $m$ (not $k$, as in [29, §19.7]), allows an elliptic integral of the second kind with $m>1$ to be recast in terms of elliptic integrals with $1/m<1$ of the first and second kinds as follows: